### v3

@@ 3.2 Enhanced FABind+
Though FABind achieves comparable docking performance with DiffDock Corso et al. (2023b) in a highly efficient way, it does not show promising advantages (docking accuracy) over existing works Corso et al. (2023b); Yan et al. (2023). By carefully investigating the evaluations of the predicted docking poses, we find the limitations of FABind can be attributed to both the inaccurate pocket/binding site predictions and the capability of the docking module. Therefore, we propose different techniques to improve both the pocket prediction and the docking modules to enhance final docking. The overall framework of our FABind+ is shown as the left part in Figure 2.
3.2.1 Pocket Prediction
Regarding pocket prediction, we found that in FABind, there are many cases where the predicted pocket position is incorrect, as shown in Figure 1. Therefore, we attempt to improve the pocket prediction method here. Specifically, we observed that FABind has proposed effective methods for accurately predicting the pocket center. However, in FABind, the overall pocket is determined by a fixed-radius sphere, which leads to many cases where the predicted pocket is not accurate. If the fixed radius is not large enough, the pocket cannot correctly contain all possible amino acids and atoms of the ligand. Since our docking network requires interactions between all atoms of the ligand and amino acids in the pocket, the size of the pocket being able to contain all atoms of the ligand is very important and can provide a good foundation for the docking module.
@@ 3.2.1 Pocket Prediction
Dynamic Pocket Radius Prediction. To enlarge the pocket coverage with the whole ligand, we propose a dynamic pocket radius prediction module. In the ideal case, even when we predict a wrong pocket center, the radius-enlarged pocket can still cover most of the atoms in the ligand. For this, we simply add a regression head for the pocket radius prediction. The label of the training data is set as the radius of each ground truth ligand conformation, which hence can be viewed as a prediction of the ligand size [^1] We also try other choices as training labels, such as the distance between the predicted pocket center and the ground truth pocket center plus the ligand radius. In practice, the ligand radius as label is best for stable training..
Denote the radius of the ground truth ligand conformation as $r$, with the updated hidden states $\mathbf{h}_{i}$ for ligand atoms from the pocket prediction module. The radius regression head, represented as $\phi_{r}$, employs a multilayer perceptron (MLP) and the loss function is the Huber loss Huber (1992):
$\hat{r}=\phi_{r}(\sum_{i}{\mathbf{h}_{i}}),\quad L_{r}=\text{Huber}(r,\hat{r}).$ (1)
Then, similar to FABind, we calculate the predicted center of the classified residues from the pocket prediction module in FABind and take it as the predicted pocket center, and the circled pocket is then in a sphere near the predicted center under the predicted radius $\hat{r}$, instead of a fixed radius. In this way, we are able to cover most of the atoms in the ligand for the further docking part.
Figure 3: Analysis of pocket prediction and the predicted ligand RMSD. “Max Distance” is the distance between the predicted pocket center and the farthest ground truth ligand atom.
3.2.2 Docking structure prediction
@@
where $\mathbb{S}$ represents the set of permutation operations applied to symmetric atoms of the molecule. In practice, we use the graph tool [^2] https://graph-tool.skewed.de to extract all permutations of a ligand graph. By incorporating permutation loss, the model directly learns to predict more rational molecular structures, streamlining the docking process and enhancing its accuracy.
Enlarge Model Size. To enhance the overall performance of the docking accuracy, we find the capability of the docking module crucial for the final result. Therefore, in this work, we simply enlarge the docking module size. However, since we still aim to maintain the efficiency of the docking speed of FABind, adding more layers will lead to slow inference. For a balance between speed and accuracy, we extend one more FABind layer for docking prediction.
3.3 A Simple Sampling Method
Recently, generative models have been popular in AI research, as well as in the biology, chemistry, and drug discovery fields Lopez et al. (2020); Bian & Xie (2021); Cheng et al. (2021). Specific to molecular docking, many works have been proposed to model the docking process in a generative way Corso et al. (2023b); Wang et al. (2023); Guo et al. (2023). The motivation behind this is to leverage the powerful sampling ability of these generative models, which can sample reasonable docking poses or pockets in the large and extensive search space. To align the line of these sampling-based methods for generative docking, we are interested in extending our FABind+ into a sampling version with minimal modifications (see the right part in Figure 2). Notably, it is obvious that for the sampling-based version, the speed has to be sacrificed, but the accuracy should be ensured and even improved.
3.3.1 Sampling method
@@
In DiffDock Corso et al. (2023b), the authors discussed the problem of the regression model due to the possibility of different pockets in a target protein, which indicates the prediction of different pocket variants is crucial. Hence, in our sampling-based FABind+, we utilize a clustering method for pocket variant prediction. Besides, a simple dropout-based conformation generation method is adopted in our model to produce conformation variants. It is important that our modified FABind+ for the sampling version requires no training of the pocket variants and conformation variant generation, which are indeed lightweight and based on the regression-based FABind+.
@@
Clustering for Pocket Variants. The pocket prediction module of FABind is designed to predict a single pocket center, which is insufficient for identifying multiple binding sites. For pocket variants, we adopt the method from Braun & Fayne (2022) to use DBSCAN for pocket clustering. This algorithm does not require the pre-specification of cluster centers and exhibits a tolerance for noise, making it a versatile choice among clustering algorithms. During each sampling process, there is a probability $p$ (with $p=0.5$) that we randomly choose a cluster from the DBSCAN output; otherwise, the initially predicted pocket center is used. The center of this cluster is then adopted as the pocket center. This strategy is aimed at maximizing diversity. Regardless of the method used to determine the pocket center, the radius predicted by the radius prediction module is employed to redefine the pocket boundaries in a spherical manner, ensuring sufficient coverage of potential ligand positions.
@@
Dropout for Conformation Generation. To sample variants of the ligand conformation, we use a very simple trick to do the sampling. Dropout Srivastava et al. (2014) is a commonly adopted method to overcome overfitting and improve generalization. It randomly drops several units in each model layer and leads to a sub-model over the full model for forward passing. In this work, we utilize the property of randomness from dropout to do the sampling. Since each sub-model produced by dropout is supervised by the ground truth conformation (label) during training, the conformation generated by the sub-model would largely be reasonable and near the correct pose. Therefore, we simply use each forward pass of the sub-model with a dropout to generate the conformation variants. The implementation is also easy, similar to Rdrop Wu et al. (2021), when generating a size of $s$ conformations for an input ligand $x$, we can simply repeat the input $x$ for $s$ times in a batch and go through one forward pass to generate $s$ conformations.
@@
Figure 4: Confidence model training pipeline. We add lightweight MLP layers upon the fixed FABind+ model for the confidence model, which incorporates a ranking loss and a classification loss for training the confidence model.
3.3.2 Confidence model
After generating multiple conformations as candidates, we need a confidence model to make the final pose selection. In this work, we introduce a lightweight confidence model with only several MLP layers after the docking module. The overall pipeline is shown in Figure 4. Specifically, the hidden states $\mathbf{h}_{i}$ before the structure/coordinates prediction are first summed and then fed into MLP layers for confidence training. During the training phase, the parameters of FABind+ are frozen, allowing for exclusive updates to the parameters of the confidence model.
We employ two loss functions here for better confidence model training. Firstly, we adopt a classification loss identical to the one used in DiffDock, which aims to predict whether the structure of a conformation w.r.t. the ground truth conformation belongs to RMSD$<$2Å. Secondly, we introduce a ranking loss that concentrates on the relative positions between conformation candidates. We use a pairwise ranking loss that is the same as the reward model training in InstructGPT Ouyang et al. (2022). Intuitively, for the conformation with a lower RMSD score, we will rank it in front of the one with a larger RMSD. These two losses have different goals, and we find they both contribute to better selection. The details are described in Appendix B.
4 Experiments
4.1 Setting
Data. We evaluate our methods on PDBbind v2020 dataset Liu et al. (2017), curated from Protein Data Bank (PDB) Burley et al. (2021). Consistent with the data split method of EquiBind, as adopted in most prior research Corso et al. (2023b); Stärk et al. (2022); Lu et al. (2022); Pei et al. (2023), we used structures published before 2019 for training and those from 2019 onwards for testing. We excluded proteins with over 1500 residues and ligands with more than 150 atoms, resulting in 17,644 samples for training, 958 for validation, and 363 for testing. Further details about our preprocessing steps are outlined in Appendix Section A.1.
Baselines. FABind+ is benchmarked against many traditional methods and deep learning models. QVINA-W, GNINA McNutt et al. (2021), SMINA Koes et al. (2013), GLIDE Friesner et al. (2004) and VINA Trott & Olson (2010) are concluded as the powerful traditional methods. Deep learning models can be categorized into two main classes: sampling-based and regression-based models. For the sampling-based models, we include DiffDock Corso et al. (2023b), and for regression-based models, we include EquiBind Stärk et al. (2022) TankBind Lu et al. (2022), E3Bind Zhang et al. (2022), and FABind Pei et al. (2023).
Evaluation Metrics. Our evaluation employs two key metrics: (1) Ligand RMSD: This metric measures the root-mean-square deviation (RMSD) between the predicted and ground-truth ligand coordinates. We report the symmetry-corrected RMSD using sPyRMSD [^3] https://github.com/RMeli/spyrmsd Meli & Biggin (2020). Detailed descriptions can be found in Appendix Section A.2. (2) Centroid Distance: This metric calculates the Euclidean distance between the centroids of the predicted and true ligand structures. It reflects binding site prediction accuracy.
Implementation Details. Initial ligand conformations are generated using the ETKDG algorithm Riniker & Landrum (2015), followed by MMFF optimization via RDKit Landrum et al. (2013), which produces the random generation of a low-energy ligand conformation. For pocket radius prediction, considering the fixed radius of FABind is 20Å, we adjust any predicted radius falling below this 20Å threshold to 20Å. This adjustment ensures the predicted pocket is at least as large as those predicted by FABind. For other hyperparameters and training configurations, we put the details in Appendix C.
@@ 4.1 Setting
Table 1: Performance of flexible blind self-docking. The number of poses that DiffDock and FABind+ sample is specified in parentheses. FABind+ without parentheses denotes regression-based performance. Methods operating solely on the CPU are marked with ”*”. All baseline results (with the exception of those from DiffDock, as reported in their paper) are sourced from Pei et al. (2023). The best results are highlighted in bold, and the second-best scores are marked with an underline.
@@
Methods 25% 50% 75% Mean 25% 50% 75% Mean Runtime (s)
traditional docking softwares
@@
deep learning-based methods
@@
FABind+ 1.2 2.6 5.8 5.2 43.5 71.1 0.4 1.0 2.9 3.5 67.5 84.0 0.16
FABind+ (10) 1.4 2.6 5.4 5.1 42.1 72.5 0.5 1.1 2.6 3.5 68.3 85.7 1.6
FABind+ (40) 1.3 2.4 5.3 5.1 43.8 73.3 0.5 1.0 2.6 3.5 69.1 86.2 6.4
4.2 Main results
@@
Blind Self-Docking Performance. Blind self-docking is the process of docking a flexible ligand to a protein without knowing the exact binding site, requiring accurate predictions of the translation, rotation, and conformation of the ligand. From the results in Table 1, our regression-based approach significantly advances the performance over the previous state-of-the-art, DiffDock. It achieves a success rate of 43.5% for ligand atomic RMSD less than 2Å, surpassing DiffDock by 5 percentage points. This model demonstrates superior accuracy across the board, as evidenced by improvements in both the mean RMSD and the percentage of predictions under 2Å and 5Å.
On the other hand, our sampling-based model shows exceptional results, especially as the sample size increases. With a sample size of 10, it performs slightly better than the regression-based model. As we expand the sample size to 40, mirroring the setting used by DiffDock, our sampling-based model delivers superior performance across most metrics. Notably, while the accuracy of highly precise predictions (RMSD <2Å) is slightly improved (43.8%), it significantly improves the prediction of more challenging targets, as indicated by an improved 75th percentile. This advancement may be credited to the implementation of clustering for multiple binding sites, which aids in correcting predictions with incorrect pocket identification.
@@
Table 2: Performance of flexible blind self-docking on unseen receptors. All baseline results are sourced from Pei et al. (2023).
@@
Methods 25% 50% 75% Mean 25% 50% 75% Mean
traditional docking softwares
@@
deep learning-based methods
@@
FABind+ 1.6 3.3 8.9 7.0 34.7 63.2 0.5 1.5 4.2 5.1 58.3 77.1
FABind+(10) 1.7 3.0 8.2 7.5 34.0 63.2 0.8 1.4 4.6 5.7 61.8 76.4
FABind+(40) 1.7 2.9 8.4 7.2 33.3 63.9 0.8 1.5 4.6 5.4 59.7 77.1
Blind Self-Docking Performance for Unseen Proteins. In this section we seek to assess the generalization capability of FABind+ on proteins not encountered during the training phase. Following previous works Zhang et al. (2022); Pei et al. (2023), we evaluate the performance of FABind+ on a set of proteins filtered based on their UniProt IDs, specifically retaining only those samples whose proteins are not seen in the training and validation stages. The results of this evaluation are summarized in Table 2. It is evident from the findings that FABind+ exhibits superior performance, surpassing all baseline methods and its predecessor, FABind, across almost all metrics. This performance underscores the effectiveness of our enhanced design in achieving robust generalization on unseen proteins.
@@
Performance Scaling with Increased Sample Size. To enhance the understanding of the sampling capacity, we depicted “Top1”, “Top5”, “Top10”, and “Perfect Selection” across varying sample sizes in Figure 5. “Perfect selection” denotes the act of selecting the sample with the lowest RMSD. DiffDock stands as a highly powerful generative model, and akin to DiffDock, our approach demonstrates significant performance gains as the sample size increases. When the sample size reaches 40, we are able to dock 51.2% of samples with an RMSD lower than 2Å. Our growth curves for the top1, top5, and top10 selections also exhibit a similar pattern to those of DiffDock, albeit at a slower pace. This deceleration may be attributed to the increased coverage of samples, where the remaining samples are considerably more challenging to dock. However, the observation of similar trends suggests that our lightweight confidence model is equally effective. In contrast to their 20M parameter model, our model is substantially smaller, with only 6M parameters and based on a simple MLP. Indeed, while our curves outperform those of DiffDock, this advantage begins with our random selection (sample size=1) outperforming DiffDock by about 15%. This does not necessarily indicate that the sampling ability brought forth by dropout or pocket clustering surpasses that of diffusion or other generative methods. However, it does imply that achieving enhanced sampling capacity in docking does not strictly require reliance on generative modeling.
Figure 5: Scaling Curve while Increasing Sample Size.
5 Analysis

### v4

@@ 3.2. Enhancing FABind to FABind+
Though FABind achieves comparable docking performance in a highly efficient manner, it does not demonstrate significant advantages in docking accuracy over existing works (Corso et al., 2023b; Yan et al., 2023). Our detailed evaluation of the predicted docking poses reveals that the limitations of FABind stem from both inaccurate pocket predictions and the capabilities of the docking module. Therefore, we propose new approaches to improve both the pocket prediction and the docking modules to enhance final docking. The overall framework of regression FABind+ is shown as the left part in Figure 2.
3.2.1. Pocket Prediction
Figure 3. Analysis of pocket prediction and predicted ligand RMSD. “Max Distance” is the distance between the predicted pocket center and the farthest ground truth ligand atom.
Regarding pocket prediction, we found that FABind was easy to predict incorrect pocket positions, resulting in bad docking pose prediction for the next stage, as illustrated in Figure 1. Therefore, we attempt to improve the pocket prediction method in this work. While FABind effectively predicts the pocket center, it uses a fixed-radius sphere to define the overall pocket, leading to inaccuracies. If the fixed radius is too small, the pocket fails to encompass all possible amino acids and ligand atoms. Since our docking network requires interactions between all ligand atoms and pocket amino acids, ensuring the pocket’s size can contain all ligand atoms is crucial for the docking module’s performance.
@@ 3.2.1. Pocket Prediction
Dynamic Pocket Radius Prediction. To enhance pocket coverage for the entire ligand, we propose a dynamic pocket radius prediction module. Ideally, even if the pocket center prediction is incorrect, the enlarged pocket radius can still encompass most ligand atoms. To achieve this, we introduce a regression head for pocket radius prediction. The labels of the training data are set as the radii of the ground truth ligands, which hence can be viewed as a prediction of the ligand size [^1] We have also tried other options as training labels, such as the distance between the predicted pocket center and the ground truth pocket center plus the ligand radius. In practice, we find that the ligand radius as the label is the best for stable training.. Based on the predicted ligand size, we add an additional buffer to account for surrounding context interactions.
Denote the radius of the ground truth ligand conformation as $r$, with the updated hidden states $\mathbf{h}_{i}$ for ligand atoms from the pocket prediction module. The radius regression head, represented as $\phi_{r}$, employs a multilayer perceptron (MLP). The goal is to minimize the Huber loss (Huber, 1992):
(1) $L_{r}=\text{Huber}(r,\hat{r}),\quad\hat{r}=\phi_{r}(\sum_{i}{\mathbf{h}_{i}}).$
Then, similar to FABind, we calculate the predicted center of the classified residues from the pocket prediction module as the predicted pocket center. The pocket is then defined as a sphere around this predicted center with the predicted radius $\hat{r}$ plus a buffer $\beta$, $\hat{r}+\beta$, instead of a fixed radius. In this way, we are able to cover most atoms in the ligand for the further docking part.
3.2.2. Docking Structure Prediction
After pocket prediction, FABind utilizes a scheduled sampling (Bengio et al., 2015) training strategy that gradually incorporates the predicted pocket for docking training. We replace this approach with teacher forcing, using the ground-truth pocket center for docking training while still applying pocket center noise for generalization. This modification stabilizes the training process. Apart from this, within the docking module, we explore various ways to enhance docking performance.
@@ 3.2.2. Docking Structure Prediction
where $\mathcal{S}$ represents the set of permutation operations applied to symmetric atoms of the molecule. In practice, we use the graph tool [^2] https://graph-tool.skewed.de to extract all permutations of a ligand graph.
3.3. From Regression to Sampling
Sampling capability is necessary for capturing multiple binding sites and conformations. Previous sampling models typically involve training a generative model for this purpose (Corso et al., 2023b; Wang et al., 2023; Guo et al., 2023). However, FABind+ demonstrates that it is not necessary to train a generative model; instead, simple techniques can be applied to enable a regression model to achieve high-quality sampling capability, with a novel light confidence model to select the final pose.
3.3.1. Sampling Method
@@
In DiffDock (Corso et al., 2023b), the authors discussed the limitations of the regression model due to the possibility of different pockets and conformations in a target protein. Hence, in our sampling-based FABind+, we utilize a clustering method for pocket variant prediction. Besides, we adopt a simple dropout-based conformation generation method (dropout sampling) to produce conformation variants. Notably, our modified FABind+ for the sampling version does not require training for pocket and conformation variant generation, making it lightweight and based on the regression-based FABind+.
@@
Clustering for Pocket Variants. The pocket prediction module of FABind is designed to predict a single pocket center, which is insufficient for identifying multiple binding sites. For pocket variants, we adopt the method from (Braun and Fayne, 2022; Yan et al., 2023) to use DBSCAN for pocket clustering. This algorithm does not require the pre-specification of cluster centers and exhibits a tolerance for noise, making it a versatile choice among clustering algorithms. During each sampling process, there is a probability $p$ (with $p=0.5$) that we randomly choose a cluster from the DBSCAN output; otherwise, the initially predicted pocket center is used. The center of this cluster is then adopted as the pocket center. This strategy aims at maximizing the diversity. Regardless of the method used to determine the pocket center, the radius predicted by the radius prediction module is employed to redefine the pocket boundaries in a spherical manner, ensuring sufficient coverage of potential ligand positions.
@@
Dropout Sampling for Conformation Generation. To sample variants of the ligand conformation, we use a very simple trick to leverage the randomness of dropout. Dropout (Srivastava et al., 2014) is a commonly adopted method to overcome overfitting and improve generalization. It randomly drops several units in each model layer and leads to a sub-model over the full model for forward passing. In this work, we utilize the property of randomness from dropout to do the sampling. During training, each sub-model produced by dropout is supervised by the ground truth conformation (label). As a result, the conformations generated by these sub-models are generally reasonable and close to the correct pose. Thus, we use each forward pass of the sub-model with dropout to generate the conformation variants. The implementation is also easy, similar to R-Drop (Wu et al., 2021), when generating a size of $s$ conformations for an input ligand x, we can simply repeat the input x for $s$ times in a batch and go through one forward pass to generate $s$ conformations.
@@
Figure 4. Confidence model training pipeline. We add lightweight MLP layers upon the fixed FABind+ model for the confidence model, which incorporates a ranking loss and a classification loss for training the confidence model.
3.3.2. Confidence Model
After generating multiple conformations as candidates, a confidence model is required to select the final pose. In this work, we introduce a lightweight confidence model comprising only several MLP layers following the sampling-based FABind+ backbonew. On one hand, this approach is advantageous over previous work (Corso et al., 2023b) as it allows for on-the-fly collection of training data. On the other hand, more complex model architectures do not necessarily yield better performance, as illustrated in Appendix C. The overall pipeline is shown in Figure 4. During the training of the sampling model, with the parameters of FABind+ frozen, we use pocket clustering and dropout sampling to explore different binding sites and conformations. The output states $\mathbf{h}_{i}$ from FABind+ are summed and then fed into the MLP layers to obtain the score, denoted as $s=f_{\theta}(\sum\mathbf{h}_{i})$. Here, $f_{\theta}$ represents the confidence model, $\mathbf{h}_{i}$ is the output embedding from FABind+, and the scalar output $s$ is utilized for confidence loss computation.
Training Objective. The training objective includes both classification loss, denoted as $L_{cls}$, and ranking loss, denoted as $L_{rank}$. The classification loss is formulated as a binary cross-entropy loss to accurately predict whether each pose has an RMSD below 2Å, following DiffDock (Corso et al., 2023b). The motivation behind incorporating ranking loss is that achieving precise 2Å classification is challenging, whereas ranking between samples is comparatively simpler. Specifically, we adopt a pairwise ranking loss inspired by InstructGPT (Ouyang et al., 2022). For each batch, each input ligand is repeated $N$ times, and there are then $N$ conformations generated after model forwarding, resulting $\binom{N}{2}$ for each protein-ligand complex. We train on all $\binom{N}{2}$ comparisons from each complex as a single batch element. According to InstructGPT, this method is computationally more efficient and reduces the likelihood of overfitting. The ranking loss is formally defined as follows:
(3) $L_{rank}=-\frac{1}{\binom{N}{2}}\mathbb{E}_{\left(\mathcal{R}_{+},\mathcal{R}_{-}\right)\sim D}\left[\log\left(\sigma\left(s_{+}-s_{-}\right)\right)\right],$
where $s$ is the scalar score predicted by the confidence model for conformation $\mathcal{R}$, $\mathcal{R}_{+}$ is the preferred conformation with lower RMSD compared to $\mathcal{R}_{-}$, and $D$ is the dataset of comparison sets.
In the implementation, each GPU batch consists of one input ligand, which is then duplicated $N=5$ times. Following model forward propagation, each sample yields $N$ predicted structures. These structures are subsequently ranked according to their computed RMSD values, which are also utilized to calculate the classification loss $L_{cls}$. The total loss for the confidence model, denoted as $L_{conf}$, is the sum of the classification loss and the ranking loss:
(4) $L_{conf}=L_{cls}+L_{rank}.$
4. Experiments
4.1. Experimental Setup
@@
Table 1. Performance of flexible blind self-docking. The number of poses that DiffDock and FABind+ sample is specified in parentheses. FABind+ without parentheses denotes regression-based performance. The results of the sampling-based FABind+ are averaged over three inference runs. Methods operating solely on CPU are marked with "*". All baseline results (with the exception of those from DiffDock, as reported in their paper) are sourced from Pei et al. (2023). The best results are highlighted in bold, and the second-best scores are marked with an underline.
@@
Method 25% 50% 75% Mean 25% 50% 75% Mean Runtime (s)
Traditional docking software
@@
Deep learning-based method
@@
Regression FABind+ 1.2   2.6   5.8 5.2 43.5 71.1 0.4 1.0 2.9   3.5 67.5 84.0   0.16
Sampling FABind+ (10) 1.3   2.7   5.4 5.2 42.4 71.6 0.5 1.1 2.8   3.5 67.8 84.6  1.6
Sampling FABind+ (40) 1.2   2.4   5.6 5.2 44.9 71.3 0.5 1.0 2.7   3.5 68.3 85.2  6.4
Dataset. We evaluate our methods on PDBbind v2020 dataset (Liu et al., 2017), curated from Protein Data Bank (PDB) (Burley et al., 2021). Consistent with the data split method of EquiBind, as adopted in most prior research (Corso et al., 2023b; Stärk et al., 2022; Lu et al., 2022; Pei et al., 2023), we used structures published before 2019 for training and those from 2019 onwards for testing. We excluded proteins with over 1500 residues and ligands with more than 150 atoms, resulting in 17,644 samples for training, 958 for validation, and 363 for testing. Further details about our preprocessing steps are outlined in Appendix A.1.
Baselines. FABind+ is benchmarked against many traditional methods and deep learning models. QVINA-W, GNINA (McNutt et al., 2021), SMINA (Koes et al., 2013), GLIDE (Friesner et al., 2004) and VINA (Trott and Olson, 2010) are concluded as the powerful traditional methods. Deep learning models can be categorized into two main classes: sampling-based and regression-based models. For the sampling-based models, we include DiffDock (Corso et al., 2023b) with different sample size, and for regression-based models, we include EquiBind (Stärk et al., 2022) TankBind (Lu et al., 2022), E3Bind (Zhang et al., 2022a), and FABind (Pei et al., 2023).
Evaluation Metrics. Our evaluation employs two key metrics: (1) Ligand RMSD, which measures the root-mean-square deviation (RMSD) between the predicted and ground-truth ligand coordinates. We report the symmetry-corrected RMSD using sPyRMSD [^3] https://github.com/RMeli/spyrmsd (Meli and Biggin, 2020). Detailed descriptions can be found in Appendix A.2. (2) Centroid Distance, which calculates the Euclidean distance between the centroids of the predicted and true ligand structures.
Implementation Details. We recognize that the capability of the docking module is essential for achieving optimal results. Ideally, larger models would trade speed for higher accuracy. However, as shown in Figure 12, we observe that larger models struggle with optimization. Based on this observation, we decided to simply extend the model by incorporating an additional FABind layer. A detailed analysis of this approach, along with other design techniques, is provided in Appendix B.
4.2. Main Results
4.2.1. Blind Self-Docking Performance.
@@ 4.2. Main Results
Blind self-docking involves docking a flexible ligand to a protein without prior knowledge of the exact binding site, requiring accurate predictions of the translation, rotation, and conformation of the ligand. As shown in Table 1, our regression-based approach significantly outperforms the powerful generative model DiffDock. It achieves a success rate of 43.5% for ligand atomic RMSD less than 2Å, surpassing DiffDock by 5.3 percentage points. This model demonstrates superior accuracy across the board, as evidenced by improvements in both the mean RMSD and the percentage of predictions under 2Å and 5Å.
On the other hand, our sampling-based model shows better results, especially as the sample size increases. As shown in Figure 5, with a sample size of 1, we achieve 35.3% of predictions under an RMSD of 2Å, which is reasonable since sampling models are less likely to produce highly accurate results with a single shot compared to the same predictive models. With a sample size of 10, the model slightly outperforms its regression-based version. As we expand the sample size to 40, similar to DiffDock’s setting, our sampling-based model delivers superior performance across most metrics. FABind+ with a sample size of 40 achieves a performance of 44.9%. In addition to accurate predictions, we maintain a fast sampling speed compared to other sampling-based models. The inference speed of the sampling-based FABind+ is 13 times faster than DiffDock under equivalent conditions. Notably, while the performance gains from the sampling version may not appear significant in the table, its strength lies in sampling multiple pockets and generating diverse conformations, which are crucial abilities that the regression model cannot achieve.
4.2.2. Blind Self-Docking Performance for Unseen Proteins.
Here we seek to assess the generalization capability of FABind+ on proteins not encountered during the training phase. Following previous works (Zhang et al., 2022a; Pei et al., 2023), we evaluate the performance of FABind+ on a set of proteins filtered based on their UniProt IDs, specifically retaining only samples whose proteins are unseen in the training and validation stages. The results of this evaluation are summarized in Table 2. It is evident from the findings that FABind+ exhibits superior performance, surpassing all baseline methods and its predecessor, FABind, across all metrics. This performance underscores the effectiveness of our enhanced design in achieving robust generalization on unseen proteins.
@@
Table 2. Performance of flexible blind self-docking on unseen receptors. Sampling-based FABind+ results are averaged over three inference runs. All baseline results are sourced from Pei et al. (2023).
@@
Method 25% 50% 75% Mean 25% 50% 75% Mean
Traditional docking software
@@
Deep learning-based method
@@
Regression FABind+ 1.6 3.3   8.9   7.0 34.7 63.2 0.5 1.5   4.2 5.1 58.3 77.1
Sampling FABind+(10) 1.6 3.2   9.0   7.4 33.3 61.8 0.6 1.4   4.3 5.7 59.0 75.0
Sampling FABind+(40) 1.6 3.3   8.8   7.1 35.4 61.1 0.6 1.5   4.9 5.3 58.3 76.3
Figure 5. Scaling curve with increasing sample size. “ Perfect Selection” refers to refers to choosing the samples with the lowest RMSD at each given sample size.
4.2.3. Performance Scaling with Increased Sample Size.
@@
To enhance the understanding of the sampling capacity, we depict “Top1”, “Top5”, “Top10”, and “Perfect Selection” across varying sample sizes in Figure 5. DiffDock stands as a highly powerful generative model, and akin to DiffDock, our approach shows significant performance gains as the sample size increases. When the sample size reaches 40, we are able to dock 51.2% of samples with an RMSD lower than 2Å. Our growth curves for the top1, top5, and top10 selections also mirror those of DiffDock. The observation of similar trends suggests that our lightweight confidence model is equally effective. In contrast to their 20M parameter model, our model is substantially smaller, with only 0.8M parameters, and based on a simple MLP. Indeed, while our curves outperform those of DiffDock, this advantage begins with our random selection (sample size = 1) outperforming DiffDock by about 15%. This does not necessarily indicate that the sampling ability brought forth by dropout or pocket clustering surpasses that of diffusion or other generative methods. However, it does imply that achieving enhanced sampling capacity in docking does not necessarily require generative training.
5. Analysis
5.1. Component Analysis